slipknot3066:

What are the solutions to this equation? (p+6)2=25 Enter your answers in the boxes. The solutions are p = or p =

dude:

\((p+6)^2=25\) Do you know how to start?

slipknot3066:

nope

slipknot3066:

lol ty

dude:

Alright what is the opposite of squaring?

slipknot3066:

Divide right?

dude:

No

dude:

Dividing is the opposite of multiplying

slipknot3066:

to square something is too multiply though

dude:

Multiplying repetitively by the same number, its not quite the same thing

slipknot3066:

so...what is it?

dude:

Doing the square root \(\sqrt{number}\)

dude:

So... \((p+6)^2=25\) Square root on \(\mathbb {both~sides}\) \(\sqrt{(p+6)^2}=\sqrt{25}\) Can you solve this?

slipknot3066:

..no

Pixel:

I can(solve that) (P+6) = 5

dude:

SInce powers and roots are opposites they cancel out \(\color{red}{\sqrt{}\color{black}{p+6}\color{red}{{^2}}}=\sqrt{25}\\ p+6=\sqrt{25}\) What is the \(\sqrt {25}\)

Pixel:

I was right :D I CAN MATH DUDER BRO

slipknot3066:

but...um there are 2 slots to chose from